frank issued 20 year bonds 2 years ago at a coupon rate of 6.14%, the bonds make semiannual payments and have a par value of $1,000, if the YTM on these bonds is 5.58% what is the current bond price?
Answer #1

Coupon = (6.14% of 1000) / 2 = 30.7

Semi annual Rate = 5.58% / 2 = 2.79%

Number of periods = 18 * 2 = 36

Current bond price = Coupon * [1 - 1 / (1 + r)^n] / r + FV / (1 + r)^n

Current bond price = 30.7 * [1 - 1 / (1 + 0.0279)^36] / 0.0279 + 1000 / (1 + 0.0279)^36

Current bond price = 30.7 * [1 - 0.371338] / 0.0279 + 371.337511

Current bond price = 30.7 * 22.532688 + 371.337511

Current bond price = $1,063.09